Find the value of x so that angleABC is a right angle, given the following coordinates : A(x,7) B(-3,x) C(-1,0)
Answer by Alan3354(6059)

Find the value of x so that angleABC is a right angle, given the following coordinates : A(x,7) B(-3,x) C(-1,0)
------------------
AB is perpendicular to BC.
slope of AB = (7-x)/(x+3)
slope of BC = x/(-3+1) = -x/2
-------------
the slopes are negative inverses
(x+3)/(7-x) = x/2
2(x+3) = x(7-x)
-x^2 + 7x = 2x + 6
x^2 - 5x + 6 = 0
x = 2
x = 3